### v1

# An X-Ray High-Frequency QPO in NGC 1365

11institutetext: College of Science, China Three Gorges University, Yichang 443002, China 22institutetext: Center for Astronomy and Space Sciences, China Three Gorges University, Yichang 443002, China 33institutetext: Purple Mountain Observatory, Chinese Academy of Sciences, Nanjing 210008, China 44institutetext: Key Laboratory of Particle Astrophysics, Institute of High Energy Physics, Chinese Academy of Sciences, 19B Yuquan Road, Beijing 100049, China
An X-Ray High-Frequency QPO in NGC 1365
Yongkang Yan 11    Peng Zhang E-mail:zhangpeng@ctgu.edu.cn11 2 2    Qingzhong Liu 33    Zhi Chang 44    Gaochao Liu 11 2 2    Jingzhi Yan 33    Xiangyun Zeng 11 2 2
(Received XXX / Accepted XXX)
This study presents the detection of a high-frequency Quasi-Periodic Oscillation (QPO) in the Seyfert galaxy NGC 1365, based on observational data obtained by the XMM-Newton in January 2004. Utilizing the Weighted Wavelet Z-transform (WWZ) and Lomb-Scargle Periodogram (LSP) methods, a QPO signal was identified at a frequency of $2.19\times 10^{-4}\ {\rm Hz}$ (4566 s), with a confidence level of 3.6 $\sigma$. The signal was notably absent in the lower 0.2 – 1.0 keV energy band, with the primary contribution emerging from the 2.0 – 10.0 keV band, where the confidence level reached 3.9 $\sigma$. Spectral analysis shows that there are multiple absorption and emission lines in the high-energy band (¿ 6 keV). The correlation between the QPO frequency ($f_{\rm QPO}$) and the mass of NGC 1365 central black hole ($M_{\rm BH}$) aligns with the established logarithmic trend observed across black holes, indicating the QPO is of high frequency. This discovery provides new clues for studying the generation mechanism of QPO in Seyfert galaxies, which helps us understand the accretion process around supermassive black holes and the characteristics of strong gravitational fields in active galactic nuclei.
Key Words.:
galaxies: active – galaxies: individual (NGC 1365) – galaxies: nuclei – X-rays: galaxies
1 Introduction
Active Galactic Nuclei (AGN) are a class of extragalactic objects characterized by intense energetic processes occurring at their cores, primarily driven by the accretion of matter onto Supermassive Black Holes (SMBHs). The presence of anisotropic structures within the galactic nuclei bestows a variety of observational characteristics across the electromagnetic spectrum (Swain et al., 2023), with the X-ray spectrum being notably shaped by the re-radiation resulting from the interaction of X-ray photons with the surrounding gas. This X-ray radiation is believed to originate from regions within the accretion disk or closer to the black hole(Ghisellini et al., 1994). A defining feature of AGN X-ray emission is its temporal variability, which can be attributed to changes in the accretion processes and the shielding effects of the environment (Lachowicz & Czerny, 2005). Quasi-periodic oscillations are a special phenomenon of some X-ray and $\gamma$-ray emission sources, which are quite common in neutron stars and black hole binaries in the Milky Way and nearby galaxies (Gierlinski et al., 2008; Pan et al., 2016). However, this phenomenon is very rare in AGN (Zhang et al., 2020), and with many early detections have been disfavored due to inadequacies in modeling the underlying broad-band noise (Alston et al., 2015).
Based on the time variability and spectral changes of radiation, an outburst cycle can be divided into five stages: quiescent state, low hard state, hard intermediate state, soft intermediate state, and high soft state (Muñoz-Darias et al., 2011). In X-ray outbursts, periodicity reappears, and its frequency is not constant. For example, in one observation of RE J1034+396, there was a change in the QPO frequency before and after (Gierlinski et al., 2008). According to the frequency size, QPO can be roughly divided into low-frequency QPO (LFQPO) and high-frequency QPO (HFQPO) (Ma et al., 2021). For black hole X-ray binary, frequency below 30 Hz is considered LFQPO (Remillard & McClintock, 2006). However, for AGN, due to their large mass, the QPO frequencies that appear are generally low, mostly at the $10^{-4}$ Hz level (Kluzniak & Abramowicz, 2002; Zhou et al., 2010; Pan et al., 2016; Zhang et al., 2018). In addition, there is a close relationship between the mass of the central black hole of AGN and the QPO frequency. Zhou et al. (2015) believes that if the $f_{\rm QPO}$ of AGN has a certain relationship with $M_{\rm BH}$, it can be considered as HFQPO. The specific relationship will be explained in the discussion section.
In a wide range of studies, QPOs often exhibit transient characteristics (Gierlinski et al., 2008; Pan et al., 2016), which is a short-lived phenomenon. There are various models to explain its origin, but a unified theory has not yet been formed (Zhang et al., 2020). The explanations for the origin of different types of QPOs also vary greatly. For LFQPO, the most commonly applicable model is the relativistic precession model (Lense & Thirring, 1918), which was first calculated by Lense and Thirring in 1918 under the conditions of ignoring viscosity and weak field approximation (since then, this precession has been called LT precession). At present, the more complete relativistic precession model is improved by Ingram et al. (2009). The origin of HFQPO may be related to the orbital motion of the inner accretion flow of the accretion disk. The frequency of HFQPO is relatively stable and does not change much with the change of luminosity. The main explanatory models are the resonance model (Abramowicz & Kluźniak, 2001), the accretion-jet instability in the accretion disk (Tagger & Pellat, 1999), and disk oscillations in different modes (g, c, p mode) (Wagoner, 1999).
Regardless of these rapidly changing mechanisms, they occur very close to black holes, and black holes have the most direct impact on them. Therefore, by studying the QPO properties of AGN, we can better understand the accretion process from a constant black hole to a supermassive black hole and the accretion disk theory around it (Zhang et al., 2020). This also provides a window for us to study the strong gravitational field around the black hole.
NGC 1365 is a large barred spiral galaxy in the Fornax galaxy cluster (Sandage & Tammann, 1983), with a distance of about 18.6 $\pm$ 0.6 Mpc (Madore et al., 1999; Silbermann et al., 1999; Springob et al., 2009; Polshaw et al., 2015; Jang et al., 2018). According to the latest calculations, its redshift value is z = 0.0051 (Koss et al., 2022). This spiral galaxy has been intensely investigated (a review of early work can be found in Lindblad (1999)). The galaxy contains an active core, and its spectral emission lines have both broad and narrow lines. The mass of the central black hole is in the range of (5 – 10) ×$\times 10^{6}\ M_{\odot}$ Maiolino & Rieke (1995), and the mass of the supermassive black hole can be derived from the emission lines in the broad-line region, in particular the broad components of hydrogen recombination lines, which can be calculated using the formula obtained in the near-infrared regionKim et al. (2010). NGC 1365 is also a typical Seyfert galaxy, and according to the Seyfert classification, it belongs to type 1.8 (Maiolino & Rieke, 1995). However, some scholars have classified it as type 1, 1.5 and 2 (e.g., Veron et al. (1980); Turner et al. (1993); Risaliti et al. (2007); Thomas et al. (2017)). This is due to the different classifications caused by the change in X-ray flux and the absorption body density along the line of sight (Hernández-García et al., 2015). The X-rays from the core include the hard continuous radiation of the active nucleus itself, the Fe-k line radiation of the rotating disk, and the thermal radiation of the surrounding stellar burst activity.
In this work, we employ the Weighted Wavelet Z-transform (WWZ) and Lomb-Scargle Periodogram (LSP) to detect and characterize a QPO signal in the X-ray observations of NGC 1365. The study’s findings provide insights into the ongoing discussion regarding the transient nature of Quasi-Periodic Oscillations (QPOs) and their potential correlation with the properties of the central black hole. The structure of this paper is as follows: Section 2 details the observations, data reduction, and analysis methods, while Section 3 summarizes the results and discusses their implications for the study of AGN and the broader astrophysics community.
2 Observations and Data Analysis
2.1 Observations and Data Reduction
Figure 1: Upper panel shows the light curve of the EPIC camera of the XMM-Newton ObsID 0205590301 in the range of 0.2 – 10.0 keV. Lower-left panel shows a two-dimensional plot of dynamic WWZ power in time-frequency space. Lower-right panel displays average WWZ power (black) and LSP value (red) curves. The blue lines represent the confidence level of 3.0, 3.5 and 4.0 $\sigma$, respectively. The orange dotted line represents 4.0 $\sigma$ confidence level calculated using the method proposed by Baluev (2008).
Figure 2: The pulse shape of the light curve is folded using period of 4566 s (two cycles are shown).
Figure 3: Part A shows WWZ power for ObsID 0205590301 at 0.2 – 1.0 keV, 1.0 – 2.0 keV, 2.0 – 6.0 keV and 6.0 – 10.0 keV. Part B, C, D and E represent the results of wavelet analysis at different energy bands, with the frequency range between two black dashed lines in Part A.
XMM-Newton is an X-ray detection satellite launched by the European Space Agency on December 10, 1999. It carries two sets of X-ray detectors, including three European Photon Imaging Cameras (EPIC; PN, MOS1, MOS2; Strüder et al. (2001); Turner et al. (2001)) and two Reflection Grating Spectrometers (2RGS;den Herder et al. (2001)). Within the span of 2003 and 2004 alone, XMM-Newton conducted 6 observations for NGC 1365. After a considerable interval, another observation was recently executed on February 21, 2024. In total, there have been 14 observations, with 10 of them being long-duration observations ($>$ 30 ks). In this paper, the observation data we used is from January 17, 2004, with an exposure time of 59.7 ks, and the observation number is ObsID 0205590301. We preprocess the observation data using the Scientific Analysis Software (SAS, version 21.0.0) provided by the XMM-Newton Science Operations Center, and extract the required scientific data using the evselect tool according to the standard procedure.
[…]
Traditional tools for studying time variability include Power Spectral Density (PSD) analysis, Short-Time Fourier Transform (STFT) of light curves and LSP. PSD and LSP are best suited for searching for strictly periodic signals, not designed for analyzing signals with time-varying frequencies. STFT can handle variable frequencies, but it largely depends on the choice of window function, and if chosen improperly, it usually produces significant side effects (Zhao et al., 2020). WWZ has a robust mapping in the frequency-time domain, especially when detecting the time evolution of parameters (such as period, amplitude, phase) describing periodic and quasi-periodic signals, WWZ shows great advantages as a periodic analysis method (Zhang et al., 2020). Examples of WWZ analysis of the evolution characteristics of QPOs in various celestial systems include Bravo et al. (2014), Ait Benkhali et al. (2020), Urquhart et al. (2022).
In this paper, a version of the WWZ modified with the Morlet mother function is utilized. (Foster, 1996). A WWZ analysis is conducted on the obtained light curve, resulting in three sets of two-dimensional color maps that depict the power spectrum over time. By observing these two-dimensional maps, we can determine whether the light curve has a period and the location of the period. At the same time, LSP method is additionally employed for further analysis and to perform comparative checks. The LSP method is different from the principle method used by WWZ. Its advantage is that it can be used to analyze the periodicity in irregular time series, by fitting the sine waves in the entire data series using the $\chi^{2}$ statistic. This method reduces the impact of irregular sampling and points out any periods or quasi-periods that may exist in the data, and it can also calculate their significance.
[…]
In addition, we also used the method proposed by (Baluev, 2008) to calculate the $4.0\ \sigma$ confidence level. Represented by the orange dotted line in Figure 1, you can see that the result is significantly more than 4.0 $\sigma$. However, this method does not take into account the impact of red noise, and for observations with a high impact of red noise, the results calculated by this method are somewhat controversial. Therefore, we can conclude that the QPO discovered in this study has a high level of confidence.
The entire time span of the EPIC light curve was folded with a period of 4566 s, and the result is shown in Figure 2. The error of each point is calculated based on the standard deviation ($68.3\ \%$) of the mean values of each phase interval. For clarity, we plotted two periods. The average count rate is about $\sim$ 3.07 counts/s. From the figure, it can be seen that the amplitude of the X-ray flux changes significantly with the phase.
In order to more accurately analyze which energy band the QPO originates from, we divided 0.2 – 10 keV into four parts: 0.2 – 1.0 keV, 1.0 – 2.0 keV, 2.0 – 6.0 keV and 6.0 – 10.0 keV. For different energy bands, we obtained the light curves using the same parameters and data processing methods. Their WWZ results(Figure 3) show that there was no QPO signal detected in 0.2 – 1.0 keV energy band, only a hint of presence at 1.0 – 2.0 keV, and the main contribution of the QPO signal comes from the 2.0 – 10.0 keV range. Similarly, to more clearly describe this difference, we calculated the confidence level for the 2.0 – 10.0 keV energy band. Through $10^{6}$ artificial light curves, we obtained a confidence level of 3.9 $\sigma$.
2.2 Time-averaged Spectra Analysis
To study the spectral properties of NGC 1365 more accurately, XSPEC (Version 12.13.1n, Arnaud (1996)) was employed for the spectral analysis. The energy range we used is data from 1.0 – 10.0 keV. As a starting point, An initial simple model, consisting of three distinct components related to AGN, was utilized: $TBabs\times(powerlaw+pexrav+apec)$ (Nardini et al., 2015), to simultaneously fit the spectra from three EPIC cameras. In this model, TBabs is the Tuebingen-Boulder ISM absorption model, representing the galactic absorption of NGC 1365, fixed at $N_{\rm H}=1.34\times 10^{20}$ cm-2 (according to Kalberla et al. (2005)). For the part reflected by the irradiated material, we used a pexrav model (Magdziarz & Zdziarski, 1995) to describe it, which has solar abundance, an inclination angle of 45, and variable intensity. It also includes a soft X-ray power law, interpreted as a dim, scattered AGN continuum. At the same time, to check whether a part of the emission line can be attributed to optically thin and collisionally ionized plasma, which is an important component of the XMM–Newton/RGS spectrum (Guainazzi et al., 2009), we added an apec thermal component with solar abundance (Smith et al., 2001). We used the above model as the initial model for fitting.
The fitting result of this model shows a significant negative deviation at 6.7 – 8.3 keV, a phenomenon that has been discussed in several previous papers (Risaliti et al., 2005, 2009). This is because there are four absorption lines in the X-ray spectrum of NGC 1365 in this range. Through spectral analysis, we identified these features as the $K\alpha$ and $K\beta$ lines of Fe xxv and Fe xxvi. Specifically, the corresponding energies of Fe xxv $K\alpha$ and Fe xxvi $K\beta$ are 6.697 keV and 6.966 keV, respectively, while those of Fe xxv $K\alpha$ and Fe xxvi $K\beta$ are 7.880 keV and 8.268 keV (Risaliti et al., 2005). Therefore, we multiplied two gabs models (Maitra et al., 2018; Klochkov et al., 2011) to describe these absorption lines.
First, we fixed the LineE of the two gabs models at 6.832 keV and 8.074 keV, respectively, and released them after the other parameters were well fitted. In the end, the LineE of the two gabs models were fitted at 6.837 keV and 7.988 keV, respectively. The new fit (Model B) brought significant improvement, with the goodness of fit is $\chi^{2}/\nu=1.27$. However, when we checked the fitting data, we found that the fitting result of the pexrav model had a large deviation, with a negative photon index power and a reflection scaling parameter very close to zero. After discussion, it was determined that the reflection model was not suitable and was therefore removed. NGC 1365 is considered to show partially obscured X-ray absorption (Risaliti et al., 2009; Maiolino et al., 2010), so we introduced the partially obscured model TBpcf (Swain et al., 2023) for description and multiplied it with the power law component.
The model is referred to as Model C: $Tbabs\times gabs\times gabs\times(apec+tbpcf\times powerlaw)$. In this model, the parameters of the gabs model first remain the same as the original ones. For the parameters of the TBpcf model, we refer to the findings of Swain et al. (2023), and fix the partial shielding fraction at 0.8. After obtaining the preliminary fitting results, we release the above parameters for further fitting, and obtain a better fit.The goodness of fit of the overall statistical data is $\chi^{2}/\nu=1.26$. After carefully observing the residual plot and logarithmic data plot, we found that there is an emission line near 6.4 keV, and the distribution of residuals near 8.0 keV is uneven. Therefore, we introduced two more gabs models and a Gaussian emission line model (Brenneman et al., 2013).
Therefore, we obtained the final fitting model(model D): $Tbabs\times gabs\times gabs\times gabs\times gabs\times(gaussian+apec+tbpcf\times powerlaw)$. Based on the original fitting model C, we first fixed the LineE of the gaussian at 6.40 keV for fitting, then released it, and finally fitted at 6.35 keV. This model provides a good fit, with an overall statistic of $\chi^{2}/\nu=1.10$. We compared the results we obtained with the fitting data from the studies of Risaliti et al. (2005), Nardini et al. (2015) and Swain et al. (2023), and the results are somewhat consistent. The best-fitting parameters and related errors are shown in Table 2. The spectrum of the best-fitting model is displayed in the middle panel of Figure 4, the top panel shows the ld plot, and the bottom panel shows the residuals.
Table 1: Best-fitting spectral parameters derived from model D.
Component Parameter Data
LineE $6.682_{-0.005}^{+0.010}$
Gabs1 Sigma $0.073_{-0.016}^{+0.018}$
Strength $0.204_{-0.018}^{+0.019}$
LineE $7.024_{-0.012}^{+0.010}$
 Gabs2 Sigma $0.022_{-0.003}^{+0.003}$
Strength $0.321_{-0.099}^{+0.166}$
LineE $7.990_{-0.053}^{+0.051}$
Gabs3 Sigma $0.262_{-0.067}^{+0.074}$
Strength $0.260_{-0.072}^{+0.087}$
LineE $9.549_{-0.289}^{+0.433}$
Gabs4 Sigma $2.092_{-0.279}^{+0.381}$
Strength $6.984_{-1.652}^{+2.721}$
LineE $6.350_{-0.011}^{+0.022}$
Gaussian Sigma (×$\times 10^{-4}$) $7.866_{-1.255}^{+3.224}$
Norm (×$\times 10^{-5}$) $2.007_{-0.431}^{+0.520}$
Apec kT $0.791_{-0.050}^{+0.044}$
Norm (×$\times 10^{-4}$) $1.470_{-0.181}^{+0.234}$
TBpcf $N_{H}$ (×$\times 10^{22}$) $7.719_{-0.395}^{+0.394}$
pcf $0.863_{-0.029}^{+0.023}$
Powerlaw $\Gamma$ $0.293_{-0.220}^{+0.191}$
Norm (×$\times 10^{-4}$) $7.231_{-1.719}^{+2.109}$
Reduced $\chi^{2}/\nu$ 1.10
Table 2: *
Notes. Errors are 90% confidence limits for one parameter. The units of the parameters in the table are the default model parameters used by XSPEC.
Figure 4: XMM-Newton spectrum of NGC 1365 (ObsID 0205590301). Top and middle panel are the logarithmic data and eeuf diagram. Black, red and blue represent the data of PN, MOS1 and MOS2, respectively. The best-fitting model of model D and the residuals ([data–model]/error) are shown in lower panel.
Figure 5: Correlation between the black hole masses and their QPO frequencies. The events reported in previous works are maked with green points and the new QPO signal detected in NGC 1365 is marked with a red square. The three lines represent the relations suggested in Remillard & McClintock (2006) and Kluzniak & Abramowicz (2002). See Zhou et al. (2010, 2015) and Zhang et al. (2017, 2018) for additional information.
3 SUMMARY AND DISCUSSION
In this paper, the observational data from the XMM-Newton X-ray telescope for the Seyfert galaxy NGC 1365, obtained in January 2004, were analyzed in this research. By employing the WWZ method to construct the time-averaged power spectrum and analyzing it with the LSP method, we identified a QPO signal with a period of 4566 s and a confidence level of 3.6 $\sigma$ throughout the observation period. However, we did not detect this signal in other XMM-Newton observational data. Therefore, we believe that this signal, like the QPO signals found in Mrk 766 (Zhang et al., 2017), 1H 0707--495 (Pan et al., 2016; Zhang et al., 2018), and Swift J1644$+$57 (Zhou et al., 2015), is a transient periodic signal.
Currently, the origin of QPO remains a topic of debate in the academic community. However, existing research generally believes that QPO signals may be generated by several mechanisms:Pulsating accretion near the Eddington limit;Instability of the inner accretion disk; X-ray hotspots orbiting the black hole or disk precession (according to the Bardeen-Petterson effect) (see Remillard & McClintock (2006); Li et al. (2003); Syunyaev (1973); Guilbert et al. (1983); Bardeen & Petterson (1975); Mukhopadhyay et al. (2003); Gangopadhyay et al. (2012). The related explanatory models include the resonance model (Abramowicz & Kluźniak, 2001), the relativistic precession model (Stella et al., 1999), the acoustic oscillation model (Rezzolla et al., 2003), the accretion-ejection instability in the accretion disk (Tagger & Pellat, 1999), and disk oscillations under different modes (g, c, p mode) (Wagoner, 1999).
At the same time, these models also predict the relationship between $f_{\rm QPO}$ and $M_{\rm BH}$. In the $f_{\rm QPO}-M_{\rm BH}$ correlation diagram, we first display some of the previous research (Remillard & McClintock, 2006; Kluzniak & Abramowicz, 2002; Zhou et al., 2010, 2015; Pan et al., 2016), As shown in Figure 5. The QPO frequencies found in RE J1034+396 (Gierlinski et al., 2008) and 1H 0707--495 (Pan et al., 2016; Zhang et al., 2018) have been considered to be HFQPOs (Zhou et al., 2010, 2015). The QPO frequency of NGC 1365 and the mass of the central black hole are plotted on the graph, demonstrating a high level of consistency. This suggests that the discovered QPO may also be classified as HFQPO. The mass data comes from the calculation obtained by Fazeli et al. (2019): the mass of the central black hole is in the range of (5 – 10) ×$\times 10^{6}\ M_{\odot}$. The origin of QPO in black hole X-ray binaries and Seyfert galaxies is still unclear, and our results can provide more information for this research.
We attempted to fit the energy spectrum ranging from 0.2 – 10.0 keV, but regardless of the model employed, we were unable to achieve a satisfactory fit. Consequently, we calculated the energy spectra for all observations of NGC 1365, and the results indicated that they exhibit minimal variance in 0.2 – 1.0 keV range. None of them exhibited QPO signals, thus leading us to believe that the processes generating QPO signals predominantly emit hard X-ray photons with energies exceeding 1.0 keV. After excluding photons without QPO signals, the signal’s confidence level was enhanced, reaching 3.9 $\sigma$. We did not observe a situation where QPOs only appeared in the high energy band in Mrk 766(Zhang et al., 2017) and 1H 0707-495(Zhang et al., 2018). This scenario likely indicates that the region where QPOs occur primarily emits high-energy X-ray radiation.
The X-ray data in the 1.0 – 10.0 keV band were fitted with the apec model, primarily for the soft band, and the TBpcf model, which accounts for partial obscuration, was used in conjunction with a powerlaw to describe the hard band. The fitting results of the spectrum show that the plasma temperature reaches 0.791 keV, which is close to the results obtained in Swain et al. (2023). Also, NGC 1365 has multiple absorption lines and emission lines in the high-energy band (¿ 6.0 keV), which has been studied many times before (Risaliti et al., 2005, 2009; Nardini et al., 2015; Guainazzi et al., 2009), and these emission lines also show puzzling asymmetry (Lena et al., 2016). At the same time, we applied the same model to perform spectral fitting again for the peak and valley regions in the light curve, and compared the model fitting results of the peaks and valleys. The analysis shows that there is no significant difference between the best fitting data of the peaks and valleys, indicating that the QPO phenomenon is not caused by the spectrum. Considering the causes of other HFQPOs, we believe that the QPO discovered in this study may also be caused by the resonance mechanism.(Abramowicz & Kluźniak, 2001)
At the same time, the gas around the black hole causes the absorption or reflection body to show strong heterogeneity, so the reflection intensity of the gas with the same column density measured by absorption is systematically stronger than expected, and there is partial obscuration (Risaliti et al., 2009; Maiolino et al., 2010), therefore, the $N_{\rm H}$ varies greatly under different observation conditions. Unlike standard Seyfert galaxies such as NGC 2617, the unstable behavior of NGC 1365 below 10.0 keV makes it difficult to obtain strictly consistent information from different observation times, which also leads to the lack of a unified conclusion on the classification of NGC 1365. The above complex properties make the spectrum of NGC 1365 difficult to describe simply. Therefore, more potential observations with high spatial resolution and high spectral resolution on NGC 1365 will help to study the sub-plasma scale environmental properties of active galaxies in depth.
Acknowledgements
This work is supported by the National Natural Science Foundation of China under grants 12203029, 12373030, 12233002, U2031205, 2021YFA0718500 and 2023AFB577. The data used in this paper were obtained from the XMM-Newton, and we express our gratitude.